OpenStudy (anonymous):

Barbara opened a store credit card to purchase a game console for $396. She put the entire purchase on the credit card. Her APR is 24.99% and the minimum payment is 5%. Assuming she makes only minimum payments, what is the remaining balance on her card after 3 months?

OpenStudy (anonymous):

If APR is 24.99%, this means that the monthly rate is 24.99/12 = 2.0825%. T0 - Time of Purchase - Principal amount is $396 T1 - 1 month after purchase - Amount owing is 396*1.020825 = $404.25 5% of that sum is $20.21 Remaining balance is 404.25-20.21 = $384.04 T2 - 2 months after purchase - Amount owing is 384.04*1.020825 = $392.04 5% of that sum is $19.60 Remaining balance after 2 months is 392.04-19.60 = $372.44 T3 - 3 months after purchase - Amount owing is 372.44*1.020825 = $380.20 5% of that sum is $19.01 Remaining balance after 3 months is 380.20 - 19.01 = $361.19 There are easier way of doing this through formulas, however its important to understand where the answer comes from rather then just sticking in the numbers into a formula. Hope this helps you! Let me know if you have any questions.